# Convergence Rates for the Maximum A Posteriori Estimator in PDE-Regression Models with Random Design

### v1

@@
[…]
Convergence Rates for the Maximum A Posteriori Estimator in PDE-Regression Models with Random Design
[…]
siebel@math.uni-heidelberg.de
Heidelberg University
{adjustwidth}
50pt50pt We consider the statistical inverse problem of recovering a parameter $\theta\in H^{\alpha}$ from data arising from the Gaussian regression problem
@@
with nonlinear forward map $\mathscr{G}:\mathbb{L}^{2}\to\mathbb{L}^{2}$, random design points $Z$ and Gaussian noise $\varepsilon$. The estimation strategy is based on a least squares approach under $\left\|\cdot\right\|_{H^{\alpha}}$-constraints. We establish the existence of a least squares estimator $\hat{\theta}$ as a maximizer for a given functional under Lipschitz-type assumptions on the forward map $\mathscr{G}$. A general concentration result is shown, which is used to prove consistency and upper bounds for the prediction error. The corresponding rates of convergence reflect not only the smoothness of the parameter of interest but also the ill-posedness of the underlying inverse problem. We apply the general model to the Darcy problem, where the recovery of an unknown coefficient function $f$ of a PDE is of interest. For this example, we also provide corresponding rates of convergence for the prediction and estimation errors. Additionally, we briefly discuss the applicability of the general model to other problems.
1   1Introduction
@@
Overview: In this paper, we consider a nonlinear Gaussian regression model with random design, i.e., for a given sample size $N\in\mathbb{N}$ and a fixed variance $\sigma\in\mathbb{R}_{>0}$, we have access to data $(Y_{i},X_{i})_{i=1}^{N}$ arising from
$Y_{i}=u_{f}(X_{i})+\sigma\varepsilon_{i},\,i=1,\dots,N,$ (1.1)
@@
where $\varepsilon_{1},\dots,\varepsilon_{N}$ are i.i.d. copies of $\varepsilon\sim N(0,1)$. Here, $X_{1},\dots,X_{N}$ are i.i.d. copies of $X$, which is drawn independently of $\varepsilon$ and follows the uniform distribution over a bounded domain $\mathcal{O}\subseteq\mathbb{R}^{d}$ with smooth boundary $\partial\mathcal{O}$. The regression function $u_{f}:\mathcal{O}\to\mathbb{R}$ denotes the solution of an (elliptic) partial differential equation (PDE), which depends nonlinearly on an unknown coefficient function $f$. The goal is to present an estimation strategy for $f$ by using the data set $(Y_{1},X_{1}),\dots,(Y_{N},X_{N})$, which is an example of a nonlinear statistical inverse problem as $f$ is not observed directly, see, e.g. Nickl [2023].
@@
Many real-world phenomena, like physical and biological laws, fit into this framework as they are often described by (non-)linear dynamical systems and partial differential equations which exhibit spatial and/or temporal dependencies. Examples range from the description of cell movements (see Tokihiro [2021]) and geophysical laws describing the atmosphere (see Kalnay [2002]), to ocean dynamics (see Bennett [2002]), turbulence (see Majda and Harlim [2012]), and fluid flows (see Cotter et al. [2009]). Moreover, this setting aligns with various frameworks from the perspective of inverse problems: examples include electrical impedance tomography (see Calderón [2006] and Isaacson et al. [2004]), photoacoustic tomography and other hybrid imaging problems (see Kuchment and Kunyansky [2015], Bal [2014], and particularly Bal and Uhlmann [2010] for a general framework), inverse scattering (see Colton and Kress [2019] and Hohage and Weidling [2015]), interacting particle models (see Nickl et al. [2024]), and reaction diffusion equations (see Nickl [2024]). In this work, we consider the so-called Darcy Problem (see, for instance Stuart [2010]), where we refer to Bonito et al. [2017] and Hohage and Pricop [2008], and the references therein, for concrete application examples. In this situation $u_{f}:\mathcal{O}\to\mathbb{R}$ describes a steady-state heat flow, i.e. the unique solution to the elliptic PDE
@@
$\begin{dcases}\hfill\operatorname{div}(f\cdot\nabla u)=g,\text{ on $\mathcal{O}$},\\ \hfill u=0,\text{ on $\partial\mathcal{O}$},\end{dcases}$ (1.2)
@@
with $\nabla$ denoting the gradient of a function and $\operatorname{div}$ the divergence operator applied to a vector field, respectively. Here $g>0$ is a heat source function on $\mathcal{O}$, which is known in advance, and $f:\mathcal{O}\to(0,\infty)$ describes an unknown conductivity (or diffusion) coefficient function. In real-world observation schemes, idealized observations that exactly match the mathematical formalism of a PDE are typically unavailable, as measurement errors introduce statistical noise, motivating the observation scheme Eq. 1.1 in this work. More precisely, the data consists of two parts - the exact solution $u_{f}$ of the Darcy problem Eq. 1.2, corrupted by the second part - the Gaussian noise. Nonetheless, the goal remains the same: to reconstruct the underlying unknown coefficient function $f$ that fits the pattern of Eq. 1.2.
@@
In the present work, we follow the approaches in Monard et al. [2020] and Nickl et al. [2020] and embed the regression model Eq. 1.1 into a more generic setting. Specifically, given a probability space $(\mathcal{Z},\mathscr{Z},\zeta)$ and a possibly nonlinear operator (forward map) $\mathscr{G}:\mathbb{L}^{2}(\mathcal{O})\to\mathbb{L}_{\zeta}^{2}(\mathcal{Z})$, we assume to have access to observations $(Y_{i},Z_{i})_{i=1}^{N}$, which arise from the equation
$Y_{i}=\mathscr{G}(\theta)(Z_{i})+\sigma\varepsilon_{i},\,i=1,\dots,N,$ (1.3)
@@
where, in this situation, $Z_{1},\dots,Z_{N}$ are i.i.d. drawn from $\zeta$. The goal is now to provide an estimation strategy for $\theta$, which defines an object in a Sobolev space $H^{\alpha}(\mathcal{O})$. Note that Eq. 1.1 corresponds to the special case where $\mathscr{G}$ is the nonlinear solution map $f\mapsto u_{f}$ of the Darcy problem. The reconstruction of parameters $\theta$ following an indirect and noisy observation scheme as in Eq. 1.3 has a long mathematical history. For an extensive overview of models with deterministic noise and a linear relationship $\theta\mapsto\mathscr{G}(\theta)$, we refer to Engl et al. [1996]. In the last twenty years a wide field has been established to also consider nonlinear and statistical inverse problems. We refer to Benning and Burger [2018], Kaltenbacher et al. [2008], Dashti et al. [2013], Dashti and Stuart [2015], Stuart [2010], Nickl [2023], Arridge et al. [2019], Bissantz et al. [2007], and Kaipio and Somersalo [2005], as well as to the references therein. For a novel proof in the Bayesian approach within this PDE-constrained regression setting, we refer to Monard et al. [2020] and Monard et al. [2021], where statistical guarantees, especially for the posterior mean, are derived under Gaussian process priors on the parameter of interest $\theta$ and analytical assumptions on the forward map $\mathscr{G}$. The applicability to PDE models is illustrated not only for the inversion of non-Abelian X-ray transforms but also for a model as in Eq. 1.1 with $u_{f}:\mathcal{O}\to\mathbb{R}$ being the unique solution of the stationary Schrödinger equation. Giordano and Nickl [2020] adopt these ideas for the Darcy problem Eq. 1.2 and provide results on posterior convergence. General assumptions to show posterior convergence are formulated in Nickl [2023]. In Nickl et al. [2020], the authors propose a (generalized) penalized least squares estimator $\hat{\theta}$ for $\theta$ in a Gaussian white noise model, which can be interpreted as minimizer of the quantity
$H^{\alpha}(\mathcal{O})\ni\theta\mapsto\left\|Y-\mathscr{G}(\theta)\right\|^{2}+\mathfrak{r}^{2}\left\|\theta\right\|^{2}_{H^{\alpha}(\mathcal{O})}\in\mathbb{R}_{\geq 0}.$
@@
They make use of M-estimation techniques to derive statistical guarantees for its prediction loss. The results are then applied for the special case of considering Eq. 1.1 with $u_{f}:\mathcal{O}\to\mathbb{R}$ being the unique solution of the Darcy problem Eq. 1.2 and also for the time-independent Schrödinger equation. In particular, they provide upper bounds for the prediction loss and show that these are in some situations minimax-optimal and coincide with rates known for linear operators. For the proof, the mathematical formulation of the observation scheme via a white noise process is very suitable, however, this scheme is rather infeasible for real-world applications. Furthermore, the authors discuss the relationship between $\hat{\theta}$ and the Bayesian approach, where $\hat{\theta}$ can also be interpreted as maximizer of the likelihood $\pi(\cdot|(Y_{i},Z_{i})_{i=1}^{N})$ of the posterior distribution $\Pi(\cdot|(Y_{i},Z_{i})_{i=1}^{N})$, under a Gaussian process priors $\Pi(\cdot)$ on $\theta$. This also results in the naming of $\hat{\theta}$ as maximum a posteriori (MAP) estimator. In contrast, we aim to prove results in the more realistic i.i.d. regression model Eq. 1.3.
For random design regression models described by Eq. 1.3, Bohr and Nickl [2023] and Nickl and Wang [2024] provide statistical and computational guarantees for a version of the maximum a posteriori estimator (MAP) $\hat{\theta}$, defined as the $D$-dimensional minimizer of the quantity
@@
$H^{\alpha}(\mathcal{O})\ni\theta\mapsto\frac{1}{N}\sum_{i=1}^{N}|Y_{i}-\mathscr{G}(\theta)(Z_{i})|^{2}+\mathfrak{r}^{2}\left\|\theta\right\|_{H^{\alpha}(\mathcal{O})}^{2}\in\mathbb{R}_{>0}.$ (1.4)
@@
$\left\|\mathscr{G}(\theta)-\mathscr{G}(\theta^{\prime})\right\|_{\mathbb{L}^{2}}\lesssim\left\|\theta-\theta^{\prime}\right\|_{\mathbb{L}^{2}}\text{ and }\left\|\mathscr{G}(\theta)-\mathscr{G}(\theta^{\prime})\right\|_{\infty}\lesssim\left\|\theta-\theta^{\prime}\right\|_{\infty}$ (1.5)
@@
uniformly over the whole parameter space. While the solution map of the time-independent Schrödinger equation meets these requirements, there is no known analogous result for the second condition for the solution map of the Darcy problem, which renders the theory of Bohr and Nickl [2023] and Nickl and Wang [2024] inapplicable. In addition, the first requirement leads to slower convergence rates, which we improve by taking the ill-posedness of the forward map $\mathscr{G}$ into account (see Nickl et al. [2020] for corresponding results in the white noise model). In section 2.2. of Nickl and Wang [2024], the authors propose conditions under which the stochastic gradient algorithm
$\vartheta_{0}:=\theta_{\operatorname{init}},\hskip 4.30554pt\vartheta_{k+1}=\vartheta_{k}+\rho\nabla\log\tilde{\pi}(\vartheta_{k}|(Y_{i},Z_{i})_{i=1}^{N})$ (1.6)
@@
converges geometrically to the minimizer $\hat{\theta}$ of Eq. 1.4 (and hence to the ground truth $\theta$ after sufficiently many iterations) in polynomial time, if the initial value $\theta_{\operatorname{init}}\in\mathbb{R}^{D}$ is chosen appropriately. We refer to section 2.2.5. (and particularly Theorem 2.8) in Nickl and Wang [2024]. The results are then applied to $\mathscr{G}$ being the solution map $f\mapsto u_{f}$ of the stationary Schrödinger equation, which meets global stability conditions Eq. 1.5. Under more general assumptions, Nickl [2023] discusses situations in which the analysis of the local concavity of the associated log-likelihood to Eq. 1.3 is sufficient to specify a stochastic gradient algorithm, for which $\hat{\theta}$ is computable in polynomial time, if the underlying parameter space is finte-dimensional. Under slightly technical changes, the methods also seem to apply to the Darcy problem. However, we refrain from discussing the computational aspects in more detail in this work due to the similarities with Nickl [2023] and Nickl and Wang [2024].
@@
A different approach of modeling real-world phenomena utilizes the description of dynamical (time-dependent) systems via stochastic partial differential equations (SPDEs), which inherently take into account the noise in the system dynamics. The SPDE framework describes uncertainty by embedding stochastic processes directly into the model equations, thereby offering an alternative for capturing the evolution of systems influenced by random fluctuations. For an overview of recent work of SPDEs we refer to Altmeyer et al. [2023], Altmeyer and Reiß [2021] and Gaudlitz and Reiß [2023] as well as the references therein. We emphasize that SPDEs are not the focus of the present work.
@@
In the present paper, we follow the approach of Monard et al. [2020], Monard et al. [2021] and Nickl et al. [2020] and consider the general nonlinear Gaussian regression model with random design (as Eq. 1.3) and define a maximum a posteriori estimator, similar to Eq. 1.4. We introduce conditions on the forward map $\mathscr{G}$, generalizing Eq. 1.5 by replacing the right-hand sides in both stability estimates with $(H^{\kappa})^{*}$- and $C^{\eta}$-norms, respectively, while also allowing polynomial growth. We present a general concentration result, where we compare the prediction loss $\mathfrak{d}^{2}(\hat{\theta},\theta)$ with some proxy $\mathfrak{d}^{2}(\tilde{\theta},\theta)$ (interpretable as approximation error) and obtain a corresponding consistency result for the MAP estimate, whenever the proxy approximates the ground truth sufficiently well. This result extends the findings of Bohr and Nickl [2023] and Nickl and Wang [2024], as we allow a generic reference function $\tilde{\theta}$, rather than only a finite-dimensional approximation of the ground truth and provide faster convergence rates for the squared prediction loss, which are now upper-bounded by $N^{-\frac{2(\alpha+\kappa)}{2(\alpha+\kappa)+d}}$, where $\kappa\in\mathbb{R}_{>0}$ indicates the degree of ill-posedness of the underlying inverse problem. Under an additional inverse stability assumption on the forward map $\mathscr{G}$, an analogous result for the estimation loss of $\hat{\theta}$ is established. For the sake of completeness, we provide an existence result for the MAP-estimator in this random design regression model by exploiting the direct method in the calculus of variations, where we adapt a corresponding proof of Nickl et al. [2020] for the white noise model. The general theory is applied on the Darcy problem. Specifically, we demonstrate that the corresponding solution map satisfies the new stability assumption with $\eta=1$. In this particular situation, we provide consistency for the MAP estimator, as well as corresponding convergence rate for prediction and estimation errors, which fills a gap in existing theory.
@@
Organisation of the paper: In Section 2 we introduce the most essential notation and the statistical model in detail. Further, we introduce and discuss the assumption made on the forward map $\mathscr{G}$. We define the MAP-estimator, for which we provide an existence result. The main results - consisting of the concentration inequality, an upper bound for the prediction error, as well as consistency for the MAP-estimator - are presented. In Section 3 we apply the main result to the special case of the Darcy problem. We provide analogously consistency for the MAP-estimator, as well as upper bounds for the prediction and estimation risk. In Appendix A, we provide detailed proofs for the results in the main part and further helping results.
Acknowledgement: This work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC-2181/1-39090098 (the Heidelberg STRUCTURES Cluster of Excellence). The author thanks Richard Nickl and Jan Johannes for their time, patience, and fruitful discussions, as well as for the visit to Richard Nickl’s research group in Cambridge in November 2023, which marked the beginning of this work.
2   2A general statistical inverse Problem

### v3

# Convergence Rates for the MAP Estimator in PDE-Regression Models with Random Design

@@
Abstract
We consider the statistical inverse problem of recovering a parameter $\theta\in H^{\alpha}$ from data arising from the Gaussian regression problem
@@ Abstract
with nonlinear forward map $\mathscr{G}:\mathbb{L}^{2}\to\mathbb{L}^{2}$, random design point $Z$ and Gaussian noise $\varepsilon$. The estimation strategy is based on a least squares approach under $\|\cdot\|_{H^{\alpha}}$-constraints. We establish the existence of a least squares estimator $\hat{\theta}$ as a maximizer for a given functional under Lipschitz-type assumptions on the forward map $\mathscr{G}$. A general concentration result is shown, which is used to prove consistency and upper bounds for the prediction error. The corresponding rates of convergence reflect not only the smoothness of the parameter of interest but also the ill-posedness of the underlying inverse problem. We apply the general model to the Darcy problem, where the recovery of an unknown coefficient function $f$ of a PDE is of interest. For this example, we also provide corresponding rates of convergence for the prediction and estimation errors. Additionally, we briefly discuss the applicability of the general model to other problems.
1 Introduction
@@
In this paper, we consider a nonlinear Gaussian regression model with random design, i.e., for a given sample size $N\in\mathbb{N}$ and a fixed variance $\sigma^{2}\in\mathbb{R}_{>0}$, we have access to data $(Y_{i},X_{i})_{i=1}^{N}$ arising from
$Y_{i}=u_{f}(X_{i})+\sigma\varepsilon_{i},\,i=1,\dots,N,$ (1)
@@
where $\varepsilon_{1},\dots,\varepsilon_{N}$ are i.i.d. copies of $\varepsilon\sim N(0,1)$. Here, $X_{1},\dots,X_{N}$ are i.i.d. copies of $X$, which is drawn independently of $\varepsilon$ and follows the uniform distribution over a bounded domain $\mathcal{O}\subseteq\mathbb{R}^{d}$ with smooth boundary $\partial\mathcal{O}$. The regression function $u_{f}:\mathcal{O}\to\mathbb{R}$ denotes the solution of an (elliptic) partial differential equation (PDE), which depends nonlinearly on an unknown coefficient function $f$. The goal is to present an estimation strategy for $f$ by using the data set $(Y_{1},X_{1}),\dots,(Y_{N},X_{N})$, which is an example of a nonlinear statistical inverse problem as $f$ is not observed directly, see, e.g. Nickl (2023).
@@
Many real-world phenomena, like physical and biological laws, fit into this framework as they are often described by (non-)linear dynamical systems and partial differential equations which exhibit spatial and/or temporal dependencies. Examples range from the description of cell movements (see Tokihiro (2021)) and geophysical laws describing the atmosphere (see Kalnay (2002)), to ocean dynamics (see Bennett (2002)), turbulence (see Majda and Harlim (2012)), and fluid flows (see Cotter et al. (2009)). Moreover, this setting aligns with various frameworks from the perspective of inverse problems: examples include electrical impedance tomography (see Calderón (2006) and Isaacson et al. (2004)), photoacoustic tomography and other hybrid imaging problems (see Kuchment and Kunyansky (2015); Bal (2014), and particularly Bal and Uhlmann (2010) for a general framework), inverse scattering (see Colton and Kress (2019) and Hohage and Weidling (2015)), interacting particle models (see Nickl et al. (2024)), and reaction diffusion equations (see Nickl (2024)). In this work, we consider the so-called Darcy Problem (see, for instance Stuart (2010)), where we refer to Bonito et al. (2017a) and Hohage and Pricop (2008), and the references therein, for concrete application examples. In this situation $u_{f}:\mathcal{O}\to\mathbb{R}$ describes a steady-state heat flow, i.e. the unique solution to the elliptic PDE
@@
$\begin{dcases}\hfill\operatorname{div}(f\cdot\nabla u)&=g,\text{ on $\mathcal{O}$},\\ \hfill u\hskip 2.84526pt&=0,\text{ on $\partial\mathcal{O}$},\end{dcases}$ (2)
@@
with $\nabla$ denoting the gradient of a function and $\operatorname{div}$ the divergence operator applied to a vector field, respectively. Here, $g>0$ is a heat source function on $\mathcal{O}$, which is known in advance, and $f:\mathcal{O}\to(0,\infty)$ describes an unknown conductivity (or diffusion) coefficient function. In real-world observation schemes, idealized observations that exactly match the mathematical formalism of a PDE are typically unavailable, as measurement errors introduce statistical noise, motivating the observation scheme eq. 1 in this work. More precisely, the data consists of two parts - the exact solution $u_{f}$ of the Darcy problem eq. 2, corrupted by the second part - the Gaussian noise. Nonetheless, the goal remains the same: to reconstruct the underlying unknown coefficient function $f$ that fits the pattern of eq. 2.
@@
In the present work, we follow the approaches in Monard et al. (2020) and Nickl et al. (2020) and embed the regression model eq. 1 into a more generic setting. Specifically, given a probability space $(\mathcal{Z},\mathscr{Z},\zeta)$ and a possibly nonlinear operator (forward map) $\mathscr{G}:\mathbb{L}^{2}(\mathcal{O})\to\mathbb{L}_{\zeta}^{2}(\mathcal{Z})$, we assume to have access to observations $(Y_{i},Z_{i})_{i=1}^{N}$, which arise from the equation
$Y_{i}=\mathscr{G}(\theta)(Z_{i})+\sigma\varepsilon_{i},\,i=1,\dots,N,$ (3)
@@
where, in this situation, $Z_{1},\dots,Z_{N}$ are i.i.d. drawn from $\zeta$. The goal is now to provide an estimation strategy for $\theta$, which defines an object in a Sobolev space $H^{\alpha}(\mathcal{O})$. Note that eq. 1 corresponds to the special case where $\mathscr{G}$ is the nonlinear solution map $f\mapsto u_{f}$ of the Darcy problem. The reconstruction of parameters $\theta$ following an indirect and noisy observation scheme as in eq. 3 has a long mathematical history. For an extensive overview of models with deterministic noise and a linear relationship $\theta\mapsto\mathscr{G}(\theta)$, we refer to Engl et al. (1996). In the last twenty years a wide field has been established to also consider nonlinear and statistical inverse problems. We refer to Benning and Burger (2018); Kaltenbacher et al. (2008); Dashti et al. (2013); Dashti and Stuart (2015); Stuart (2010); Nickl (2023); Arridge et al. (2019); Bissantz et al. (2007), and Kaipio and Somersalo (2005), as well as to the references therein. For a novel proof in the Bayesian approach within this PDE-constrained regression setting, we refer to Monard et al. (2020) and Monard et al. (2021), where statistical guarantees, especially for the posterior mean, are derived under Gaussian process priors on the parameter of interest $\theta$ and analytical assumptions on the forward map $\mathscr{G}$. The applicability to PDE models is illustrated not only for the inversion of non-Abelian X-ray transforms but also for a model as in eq. 1 with $u_{f}:\mathcal{O}\to\mathbb{R}$ being the unique solution of the stationary Schrödinger equation. Giordano and Nickl (2020) adopt these ideas for the Darcy problem eq. 2 and provide results on posterior convergence. General assumptions to show posterior convergence are formulated in Nickl (2023). In Nickl et al. (2020), the authors propose a (generalized) penalized least squares estimator $\hat{\theta}$ for $\theta$ in a Gaussian white noise model, which can be interpreted as minimizer of the quantity
$H^{\alpha}(\mathcal{O})\ni\theta\mapsto\|Y-\mathscr{G}(\theta)\|^{2}+\mathfrak{r}^{2}\|\theta\|^{2}_{H^{\alpha}(\mathcal{O})}\in\mathbb{R}_{\geq 0}.$
@@
They make use of M-estimation techniques to derive statistical guarantees for its prediction loss. The results are then applied for the special case of considering eq. 1 with $u_{f}:\mathcal{O}\to\mathbb{R}$ being the unique solution of the Darcy problem eq. 2 and also for the time-independent Schrödinger equation. In particular, they provide upper bounds for the prediction loss and show that these are in some situations minimax-optimal and coincide with rates known for linear operators. While the white noise model provides a mathematically convenient formulation for theoretical analysis, particularly in deriving statistical guarantees, it is often infeasible in real-world applications. More realistic observation schemes involve discrete measurements at deterministic design points, which can be chosen, for instance, to optimally cover the domain under certain regularity assumptions (see Daon (2025) and Teckentrup (2020) for instance). In our framework, we instead consider a random design regression model, which is analytically convenient and aligns well with established approaches in nonparametric statistical theory: For random design regression models described by eq. 3, Bohr and Nickl (2023) and Nickl and Wang (2024) provide statistical and computational guarantees for a Tikhonov-regularized estimator, defined as the $D$-dimensional minimizer of the quantity
@@
$H^{\alpha}(\mathcal{O})\ni\theta\mapsto\frac{1}{N}\sum_{i=1}^{N}|Y_{i}-\mathscr{G}(\theta)(Z_{i})|^{2}+\mathfrak{r}^{2}\|\theta\|_{H^{\alpha}(\mathcal{O})}^{2}\in\mathbb{R}_{>0}.$ (4)
@@
$\|\mathscr{G}(\theta)-\mathscr{G}(\theta^{\prime})\|_{\mathbb{L}^{2}}\lesssim\|\theta-\theta^{\prime}\|_{\mathbb{L}^{2}}\text{ and }\|\mathscr{G}(\theta)-\mathscr{G}(\theta^{\prime})\|_{\infty}\lesssim\|\theta-\theta^{\prime}\|_{\infty}$ (5)
@@
uniformly over the whole parameter space. While the solution map of the time-independent Schrödinger equation meets these requirements, there is no known analogous result for the second condition for the solution map of the Darcy problem, which renders the theory of Bohr and Nickl (2023) and Nickl and Wang (2024) inapplicable. In addition, the first requirement leads to slower convergence rates, which we improve by taking the ill-posedness of the forward map $\mathscr{G}$ into account (see Nickl et al. (2020) for corresponding results in the white noise model). Furthermore, the authors discuss the relationship between $\hat{\theta}$ and the Bayesian approach for inverse problems, where $\hat{\theta}$ can also be interpreted as maximizer of the likelihood $\pi(\cdot|(Y_{i},Z_{i})_{i=1}^{N})$ of the posterior distribution $\Pi(\cdot|(Y_{i},Z_{i})_{i=1}^{N})$, under a Gaussian process prior $\Pi(\cdot)$ on $\theta$. This also results in the naming of $\hat{\theta}$ as maximum a posteriori (MAP) estimator. In section 2.2. of Nickl and Wang (2024), the authors propose conditions under which the stochastic gradient algorithm
$\vartheta_{0}:=\theta_{\operatorname{init}},\hskip 4.30554pt\vartheta_{k+1}=\vartheta_{k}+\rho\nabla\log\tilde{\pi}(\vartheta_{k}|(Y_{i},Z_{i})_{i=1}^{N})$ (6)
@@
converges geometrically to the minimizer $\hat{\theta}$ of eq. 4 (and hence to the ground truth $\theta$ after sufficiently many iterations) in polynomial time, if the initial value $\theta_{\operatorname{init}}\in\mathbb{R}^{D}$ is chosen appropriately. We refer to section 2.2.5. (and particularly Theorem 2.8) in Nickl and Wang (2024). The results are then applied to $\mathscr{G}$ being the solution map $f\mapsto u_{f}$ of the stationary Schrödinger equation, which meets global stability conditions eq. 5. Under more general assumptions, Nickl (2023) discusses situations in which the analysis of the local concavity of the associated log-likelihood to eq. 3 is sufficient to specify a stochastic gradient algorithm, for which $\hat{\theta}$ is computable in polynomial time, if the underlying parameter space is finte-dimensional. Building on the proof strategy of Nickl and Wang (2024), the authors of Altmeyer (2024) establish nonasymptotic sampling guarantees for a gradient-based MCMC algorithm relative to a surrogate posterior in high-dimensional generalized linear models (see Section 2.4), a framework that also encompasses the Darcy problem discussed in Section 3.4. Hence, we refrain from discussing the computational aspects in more detail in this work due to the similarities with Nickl (2023); Nickl and Wang (2024) and Altmeyer (2024).
@@
A different approach of modeling real-world phenomena utilizes the description of dynamical (time-dependent) systems via stochastic partial differential equations (SPDEs), which inherently take into account the noise in the system dynamics. The SPDE framework describes uncertainty by embedding stochastic processes directly into the model equations, thereby offering an alternative for capturing the evolution of systems influenced by random fluctuations. For an overview of recent work of SPDEs we refer to Altmeyer et al. (2023); Altmeyer and Reiß (2021) and Gaudlitz and Reiß (2023) as well as the references therein. We emphasize that SPDEs are not the focus of the present work.
@@
In the present paper, we follow the approach of Monard et al. (2020, 2021) and Nickl et al. (2020) and consider the general nonlinear Gaussian regression model with random design (as eq. 3) and define a maximum a posteriori estimator, similar to eq. 4. We introduce conditions on the forward map $\mathscr{G}$, generalizing eq. 5 by replacing the right-hand sides in both stability estimates with $(H^{\kappa})^{*}$- and $C^{\eta}$-norms, respectively, while also allowing polynomial growth. We present a general concentration result, where we compare the prediction loss $\mathfrak{d}^{2}(\hat{\theta},\theta)$ with some proxy $\mathfrak{d}^{2}(\tilde{\theta},\theta)$ (interpretable as approximation error) and obtain a corresponding consistency result for the MAP estimate, whenever the proxy approximates the ground truth sufficiently well. This result extends the findings of Bohr and Nickl (2023) and Nickl and Wang (2024), as we allow a generic reference function $\tilde{\theta}$, rather than only a finite-dimensional approximation of the ground truth and provide faster convergence rates for the squared prediction loss, which are now upper-bounded by $N^{-\frac{2(\alpha+\kappa)}{2(\alpha+\kappa)+d}}$, where $\kappa\in\mathbb{R}_{>0}$ indicates the degree of ill-posedness of the underlying inverse problem. Under an additional inverse stability assumption on the forward map $\mathscr{G}$, an analogous result for the estimation loss of $\hat{\theta}$ is established. For the sake of completeness, we provide an existence result for the MAP-estimator in this random design regression model by exploiting the direct method in the calculus of variations, where we adapt a corresponding proof of Nickl et al. (2020) for the white noise model. The general theory is applied on the Darcy problem. Specifically, we demonstrate that the corresponding solution map satisfies the new stability assumption with $\eta=1$. In this particular situation, we provide consistency for the MAP estimator, as well as corresponding convergence rate for prediction and estimation errors, which fills a gap in existing theory.
x
@@
Organisation of the paper: In section 2 we introduce the most essential notation and the statistical model in detail. Further, we introduce and discuss the assumption made on the forward map $\mathscr{G}$. We define the MAP-estimator, for which we provide an existence result. The main results - consisting of the concentration inequality, an upper bound for the prediction error, as well as consistency for the MAP-estimator - are presented. In section 3 we apply the main result to the special case of the Darcy problem. We provide analogously consistency for the MAP-estimator, as well as upper bounds for the prediction and estimation risk. In section 4, we provide detailed proofs for the results in the main part and further helping results.
2 A general statistical inverse Problem